OpenStudy (anonymous):

20k + 50 Factor the expression and please tell me how you did it. This is not homework this is a fun thing i like to do and since it is veterans day and i have no school i am doing it. I would like the question answered by sombody who is not Mehek14.

Mehek (mehek14):

first find the GCF of both numbers

OpenStudy (anonymous):

10

Mehek (mehek14):

so 10 will be on the outside \(\tt{10(?????)}\)

Mehek (mehek14):

divide 20k/10 and 50/10 = ?

OpenStudy (anonymous):

I DON'T KNOW WHAT K IS.

Mehek (mehek14):

you don't need to know what k is

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Mehek (mehek14):

leave k then ._. divide 20/10 if that helps -.-

OpenStudy (anonymous):

2

OpenStudy (anonymous):

and 5

Mehek (mehek14):

so 20k/10 = 2k and 50/10 = 5 put that in the parenthesis \(\tt{10(2x+5)}\)

Mehek (mehek14):

and then if you multiply to check : 10*2k = 20 10 * 5 = 50 20k + 50 = 20k + 50